Calculate inflation between 2014-2015 and 2015-2016 based on the following CPI data.

YEAR

CPI

2014

215.303

2015

214.537

2016

218.056

2014-2015: _____________________

2015-2016: _____________________

What do you call what happened between 2015 and 2016? ________________________

Calculate the unemployment for each of the years shown based on the following data:

YEAR

Employed

Labor Force

2013

145,362,000

154,287,000

2014

139,877,000

154,142,000

2015

139,064,000

153,889,000

2016

139,869,000

153,617,000

2013:   _______________

2014:   _______________

2015:   _______________

2016:   _______________
Answer #1

inflation between 2014-2015=214.537/215.303-1=-0.36%

inflation between 2015-2016=218.056/214.537-1=1.64%

What do you call what happened between 2015 and 2016=price levels increased.